A 60 Ω resistor is made by winding platinum wire into a coil at 20°C. If the platinum has resistivity

r=10.4x10-8 Wm at 20°C and the platinum wire has a diameter of 0.10 mm, what length of wire is

needed?


1
Expert's answer
2021-10-25T17:52:09-0400

We use the formula

"R=\frac{\rho*L}{A}"

Where R= resistance in ohms

"\rho" = resistivity in Wm

L = Length of wire

A = Cross sectional area

We have,

R=6"\Omega", "\rho" =10.4 x 10-8Wm at 20°c

radius (r)= "\frac{diameter}{2}" = "\frac{0.10mm}{2}" = 0.05mm

We convert to meters

0.05mm = 5 x 10-5m

Now, we find the Area

Area = "\pi"r2

Where "\pi" = "\frac{22}{7}", r = 5 x 10-5

Area = "\frac{22}{7}" * (5 x 10-5)2

= "\frac{22}{7}" * 2.5 x 10-9

= 7.86 x 10-9

Now, using the formula

"R=\frac{\rho*L}{A}"

Making L the subject of the formula

"L=\frac{R*A}{\rho}"

= "\frac{60 *7.86*10^{-9}}{10.4 * 10^{-8}}"

= "\frac{4.716*10^{-7}}{10.4*10^-{-8}}"

= "\frac{1179}{260} = 4.53m (2 d.p)"

Therefore, the length L= 4.53m